Sandra would like to organize BAL as either an LLC (taxed as a sole proprietorship) or a C corporation. In either form, the entity is expected to generate an 15 percent annual before-tax return on a $680,000 investment. Sandra’s marginal income tax rate is 37 percent and her tax rate on dividends and capital gains is 23.8 percent (including the 3.8 percent net investment income tax). If Sandra organizes BAL as an LLC, she will be required to pay an additional 2.9 percent for self-employment tax and an additional 0.9 percent for the additional Medicare tax. BAL’s income is not qualified business income (QBI) so Sandra is not allowed to claim the QBI deduction. Assume that BAL will distribute all of its after-tax earnings every year as a dividend if it is formed as a C corporation. (Round your intermediate computations to the nearest whole dollar amount.)

a. How much cash after taxes would Sandra receive from her investment in the first year if BAL is organized as either an LLC or a C corporation?
Answer #1

a. Cash after taxes -

1) If BAL is organized as LLC

Annual before tax income :- 680000* 15% = $102000

Income tax = 102000*37% + 102000* 2.9% + 102000*0.9%

= 37740+2958+918 = $41616

Cash after tax= 102000-41616 =$ 60834

2) If BAL is organized as C corporation-

Annual before tax income - $102000

Income tax on company's earrings - 102000*21% =$21420

Income after tax = 102000-21420= $80580

Dividend received = $80580

Tax on dividend received in hands of Sandra= 80580*23.8%= $19178

Income after tax = 80580-19178 = $ 61402